# EarthMarker: A Visual Prompt Learning Framework for Region-level and Point-level Remote Sensing Imagery Comprehension

### v1

@@ V-E Region Captioning
For the brief region captioning, the test set of DIOR-RSVG [19] is employed. Specifically, we adopt boxes as the visual prompt and a text prompt, such as “Please provide a brief caption of each marked region in the image,” to prompt EarthMarker to concisely describe the content of the specified region using a brief caption. Similar to the image captioning task, metrics like BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, and SPICE are used to evaluate EarthMarker and other MLLMs and visual promoting models in region target understanding.
As displayed in Tab. IV, on the DIOR-RSVG test set, compared to other MLLMs such as Qwen-VL-Chat, GeoChat, SPhinx, EarthGPT, and visual prompting models such as Sphinx-V, ViP-LLava, and GLAMM, EarthMarker shows improvements of 7.72, 9.99, 11.38, 12.67, blue12.49, 7.92, 10.97, 165.67, and 17.50 in BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, CIDER and SPICE, respectively. Furthermore, we visualize the instances of EarthMarker and other models on the region captioning task, as shown in Fig. 4. In complex RS scenarios with numerous targets and extensive geographic coverage, EarthMarker can accurately identify and describe various specified targets, which are challenging tasks for other models.
@@ V-F Complex Visual Reasoning
Specifically, the response generated by GPT-4V does not specify the exact categories of each marked region. Additionally, GPT4V incorrectly describes Region 1 as “Region 1 supports all other regions indirectly by housing vehicles and equipment necessary for ground operations, including baggage handling, maintenance, and possibly emergency services”, whereas Region 1 actually contains fuel storage tanks. The Vip-LLava model incorrectly identifies all regions as containing airplanes. Note that Sphinx-V correctly identifies the types of objects in each region and analyzes the internal relationship of some regions. For example, the Sphinx-V answer “they are positioned in close proximity to each other, suggesting they are within the same area of the airport. The tanks are similar in appearance and likely serve the same function.” However, it fails to provide a comprehensive analysis of the relationships among all four regions, and incorrectly states that the airplane in Region 4 is in motion.
@@ VI Conclusion
In this paper, the fine-grained MLLM called EarthMarker, the first visual prompting model specifically designed for the RS domain, is proposed. Moreover, the RS visual prompting instruction dataset called RSVP is constructed for the first time, facilitating the development of fine-grained RS imagery comprehension. Furthermore, the visual prompts learning framework is developed. In particular, the shared visual encoding method is developed to uniformly refine multi-scale visual features and visual prompt content, which is beneficial for comprehensively understanding the interplay between visual prompts and the holistic image. Subsequently, the referring areas in the input are replaced by the proposed hybrid representation before being fed into the LLM to instruct the model to comprehend referring areas and toward specific predictions. Employing the RSVP-3M and the visual prompt learning framework, EarthMarker is equipped with multi-granularity visual understanding capability at the image, region, and point levels, making it able to simultaneously perform comprehensive and intelligent analysis in real-world scenarios. In the future, we plan to incorporate a broader range of visual modalities into EarthMarker, enhancing its multi-source imagery comprehension capabilities. In addition, we plan to support free-form shapes as visual marks to adjust the referring granularity flexibly.

### v2

@@ V-E Region Captioning
For the brief region captioning, the test set of DIOR-RSVG [19] is employed. Specifically, we adopt boxes as the visual prompt and a text prompt, such as “Please provide a brief caption of each marked region in the image,” to prompt EarthMarker to concisely describe the content of the specified region using a brief caption. Similar to the image captioning task, metrics like BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, and SPICE are used to evaluate EarthMarker and other MLLMs, visual promoting models in region target understanding.
As displayed in Tab. IV, on the DIOR-RSVG test set, compared to other MLLMs such as Qwen-VL-Chat, GeoChat, SPhinx, EarthGPT, and visual prompting models such as Sphinx-V, ViP-LLava, and GLAMM, EarthMarker shows improvements of 7.72, 9.99, 11.38, 12.67, blue12.49, 7.92, 10.97, 165.67, and 17.50 in BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, CIDER and SPICE, respectively. Furthermore, we visualize the instances of EarthMarker and other models on the region captioning task, as shown in Fig. 4. In complex RS scenarios with numerous targets and extensive geographic coverage, EarthMarker can accurately identify and describe various specified targets, which are challenging tasks for other models. For example, compared to the proposed EarthMarker, the current popular MLLMs like GeoChat and EarthGPT are still inferior in regional comprehension. This excellent performance is mainly attributed to the region-text training and the dual instructions design of visual prompts and language.
@@ V-F Complex Visual Reasoning
Specifically, the response generated by GPT-4V does not specify the exact categories of each marked region. Additionally, GPT4V incorrectly describes Region 1 as “Region 1 supports all other regions indirectly by housing vehicles and equipment necessary for ground operations, including baggage handling, maintenance, and possibly emergency services”, whereas Region 1 actually contains fuel storage tanks. The Vip-LLava model incorrectly identifies all regions as airplanes. Note that Sphinx-V correctly identifies the types of objects in each region and analyzes the internal relationship of some regions. For example, the Sphinx-V answer “They are positioned in close proximity to each other, suggesting they are within the same area of the airport. The tanks are similar in appearance and likely serve the same function.” However, it fails to provide a comprehensive analysis of the relationships among all four regions, and incorrectly states that the airplane in Region 4 is in motion.
@@ VI Conclusion
In this paper, the fine-grained MLLM called EarthMarker, the first visual prompting model specifically designed for the RS domain, is proposed. Furthermore, the RS visual prompting instruction dataset called RSVP is constructed for the first time, facilitating the development of fine-grained RS imagery comprehension. In addition, the visual prompts learning framework is developed. Particularly, the shared visual encoding method is developed to uniformly refine multi-scale visual features and visual prompt content, which is beneficial for comprehensively understanding the interplay between visual prompts and the holistic image. Subsequently, the referring areas in the input are replaced by the proposed hybrid representation before being fed into the LLM to instruct the model to comprehend referring areas and toward specific predictions. Employing the RSVP-3M and the visual prompt learning framework, EarthMarker is equipped with multi-granularity visual understanding capability at the image, region, and point levels, making it able to simultaneously perform comprehensive and intelligent analysis in real-world scenarios. In the future, we plan to incorporate a broader range of visual modalities into EarthMarker, enhancing its multi-source imagery comprehension capabilities. In addition, we plan to support free-form shapes as visual marks to adjust the referring granularity flexibly.